# A New Algorithm for Whitney Stratification of Varieties

### v1

@@ 14 Return $W_{\bullet}$
The termination of the algorithm is clear. Note that the singular locus of any variety consists of the the union of the singular loci of its $\mathbb{Q}$-irreducible components together with the intersections of all its $\mathbb{Q}$-irreducible components. Thus lines 4 and 6 in Whitney guarantee the smoothness of all strata of the output of $\text{\bf Whitney}(X)$. Further, the fact that Whitney’s condition (B) is satisfied for all pair of the strata of the output of $\text{\bf Whitney}(X)$ follows immediately from Theorem 3.6. ∎
@@
3.3. Minimizing a Given Whitney Stratification
Our first goal here is to show how to compute, for a given $X\subset\mathbb{C}^{n}$ and a random point $y$ in a subvariety $Y\subset X$, the sequence $m_{\bullet}(X,y)$ without being given $y$ explicitly. Next, we will show that the sequence $m_{\bullet}(X,\bullet)$ is constant on a Zariski-open subset of $Y$, if $Y$ is $\mathbb{Q}$-irreducible. These two things combined allow us to compute, probabilistically, the “generic” multiplicities $m_{\bullet}(X,y)$, $y\in Y$. Combining this with Theorem 2.15 we obtain a procedure to produce the unique minimal Whitney stratification of a complex algebraic variety given a Whitney stratification computed by Whitney.
Suppose for the moment that $Y=\mathbf{V}(f_{1},\dots,f_{c})$ where $c\leq n$ and $f_{1},\dots,f_{c}$ is such that $F:=(f_{1},\dots,f_{c},\ell_{c+1},\dots,\ell_{n})$ is a reduced regular sequence, where the $\ell_{i}$ are generic degree one polynomials in $\mathbf{x}$, i.e. the determinant of $\text{\bf J}(F)$ is non-zero at every point of $Y$. Let now
ϕ::italic-$\phi:$ $\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$
$p:=(p_{1},\dots,p_{n})\mapsto(f_{1}(p),\dots,f_{c}(p),\ell_{c+1}(p),\dots,\ell_{n}(p)),$
and let $\tilde{X}$ and $\tilde{Y}$ be the closures of the images under $\phi$ of $X$ and $Y$ respectively. Denote by $z_{1},\dots,z_{n}$ the coordinates on the codomain of $\phi$. Note that the origin $0\in\mathbb{C}^{n}$ lies in $\tilde{Y}$. Now we can compute the polar varieties $\delta_{i}(\tilde{X},0)=\delta_{i}(\tilde{X})$ and their multiplicities at 00 by choosing linear spaces through the origin. This gives us the correct multiplicities because of the following fact.
@@ 3.3. Minimizing a Given Whitney Stratification
Since $F$ is a reduced regular sequence, $\phi$ defines a local isomorphism near any $y\in Y$ by the inverse function theorem. This means that for any $y\in Y$, $\phi$ induces an isomorphism of completions
$\phi^{*}_{y}:\widehat{\mathscr{O}}_{\mathbb{C}^{n},\phi(y)}\rightarrow\widehat{\mathscr{O}}_{\mathbb{C}^{n},y}.$
Now, since the Hilbert-Samuel multiplicity of a local ring $(R,\mathfrak{m})$ depends only on $\operatorname{gr}_{\mathfrak{m}}(R)$ and since $\operatorname{gr}_{\mathfrak{m}}(R)=\operatorname{gr}_{\mathfrak{m}}(\widehat{R})$, the result follows. ∎
Next suppose that $Y\subset X$ is equidimensional of codimension $c$ but no longer necessarily cut out by a reduced regular sequence. Choose, at random, a sequence $f_{1},\dots,f_{c}\in I(Y)$, for example as random linear combinations of given generators of $I(Y)$. Defining, $Z:=\mathbf{V}(f_{1},\dots,f_{c})$ we then have the following
@@
Use the notations above. If $f_{1},\dots,f_{c}$ are sufficiently generic, there exists a Zariski-open subset $U\subset\mathbb{C}^{n}$ such that
$Y\cap U=Z\cap U$
and $Y\cap U$ is dense in $Y$. Moreover, for sufficiently generic degree one polynomials $\ell_{c+1},\dots,\ell_{n}$, the sequence
$F:=(f_{1},\dots,f_{c},\ell_{c+1},\dots,\ell_{n})$
is a reduced regular sequence on $Y$.
@@
Since $f_{1},\dots,f_{c}$ is sufficiently generic, $f_{1},\dots,f_{c}$ defines a maximal regular sequence in $I(Y)$. Hence there exists some polynomial $h$ such that $hI(Y)\subset\langle f_{1},\dots,f_{c}\rangle$. Choosing $U:=\mathbb{C}^{n}-\mathbf{V}(h)$ proves the first part of the statement. Now, at every point $p\in U$, we have $I(Y)\mathbb{C}[\mathbf{x}]_{I(p)}=\langle f_{1},\dots,f_{c}\rangle\mathbb{C}[\mathbf{x}]_{I(p)}$. This shows that $f_{1},\dots,f_{c}$ defines a radical ideal at every point $p\in U$. By Exercise 12.11 in [3], the same remains true for the sequence $f_{1},\dots,f_{c},\ell_{c+1},\dots,\ell_{n}$ where the $\ell_{i}$ are sufficiently generic degree one polynomials in $\mathbf{x}$, proving the proposition. ∎
@@
Let $m_{\bullet}(\tilde{X},0)$ be the multiplicity sequence constructed as above, with $Y$ replaced by $Z$. Then $m_{\bullet}(\tilde{X},0)=m_{\bullet}(X,y)$ for every $y\in\phi^{-1}(0)$.
@@
Choose $U\subset\mathbb{C}^{n}$ as in Proposition 3.11. By the same proposition, with probability $1$, the intersection $Y\cap\mathbf{V}(\ell_{c+1},\dots,\ell_{n})\cap U\subset\phi^{-1}(0)$ is not empty. For any choice of $y\in Y\cap\mathbf{V}(\ell_{c+1},\dots,\ell_{n})\cap U$ we then have $m_{\bullet}(\tilde{X},0)=m_{\bullet}(X,y)$ by Proposition 3.10. ∎
@@
5 ϕitalic-$\phi\leftarrow$ the map given by $\mathbf{x}\mapsto(f_{1},\dots,f_{c},\ell_{c+1},\dots,\ell_{n})$
6 Compute ideals defining $\phi(X)=\delta_{0}(\phi(X),0),\dots,\delta_{\dim(X)-1}(\phi(X),0)$
7 Using these ideals, Return $(m_{0}(\delta_{0}(\phi(X),0)),\dots,m_{0}(\delta_{\dim(X)-1}(\phi(X),0)))$
@@ 6 Compute ideals defining $\phi(X)=\delta_{0}(\phi(X),0),\dots,\delta_{\dim(X)-1}(\phi(X),0)$
In line 6, an ideal defining $\phi(X)$ is again computed using Gröbner basis computations with block orders, see e.g. Proposition 15.30 in [3]. Using the notation of Proposition 3.3, ideals defining the necessary local polar varieties can then be computed as ideals of the form
@@
In line 7, to compute the required multiplicities we use degree computations via Gröbner basis as described in [6] and implemented in the SegreClasses Macaulay2 package [5], see in particular [6, Theorem 5.3]. Alternatively one can use standard basis computations, see e.g. [12] for details.
Our next goal is the following: Suppose that $Y\subset X$ is $\mathbb{Q}$-irreducible. Then we want to show that the sequence $m_{\bullet}(X,\bullet)$ is constant on a Zariski-dense open subset of $Y$. This is concluded as follows from Theorem 2.15:
@@
Given a variety $X$ and $\mathbb{Q}$-irreducible $Y\subset X$, there exists a Zariski-dense open subset $U\subset Y$ such that $m_{\bullet}(X,\bullet)$ is constant on $U$.
@@
Suppose $\dim(Y)=e$. Using Proposition 3.11 and the same argument as in the proof of Proposition 3.10 we may suppose that $Y=\mathbf{V}(x_{e+1},\dots,x_{n})$ since we are computing local multiplicities and they are preserved by the map $\phi$ used in the proof of Proposition 3.10. Now, the set of points $W\subset Y$ at which the pair $(X,Y)$ fails to satisfy Whitney’s condition (B) is Zariski-closed in $Y$ and defined over $\mathbb{Q}$ (this can be seen for example by using Remark 4.2 in [4]). Let $U:=Y-W$. By Theorem 2.15, every point $y\in U$ has a euclidean neighborhood on which the sequence $m_{\bullet}(X,\bullet)$ is equal to $m_{\bullet}(X,y)$. But $U$ is connected, therefore $m_{\bullet}(X,\bullet)$ is constant on $U$, proving the theorem. ∎
@@ 7 Return $W_{\bullet}$
The termination is clear. Suppose that we have a triple of varieties $W\subset Z\subset Y$ with $W$ and $Z$ irreducible and that for generic $w\in W$ and $z\in Z$ we have $m_{\bullet}(Y,w)=m_{\bullet}(Y,z)$. Then there exists a Zariski-dense subset $U$ of $Z$ such that $U\cap W\neq\emptyset$ and such that $m_{\bullet}(Y,\bullet)$ is constant on $U$. Hence the points in $W$ at which the pair $(Y,Z)$ does not satisfy Whitney’s condition (B) have at least codimension one in $W$ by Theorem 2.15. In the situation of WhitneyMinimize, $W$ can thus be removed without destroying the property of the output flag $W_{\bullet}$ being a Whitney stratification. This proves the correctness. ∎
@@
Note that we may also combine WhitneyMinimize directly with WhitPoints: For given $X$ and irreducible $Y$ let $h$ be the output of WhitPoints($X,Y$) and let $A:=Y\cap\mathbf{V}(h)$. We know that the set of points in $Y$ where Whitney’s condition (B) fails to hold with respect to $X$ and $Y$ is contained in $A$. We may then compute the irreducible components $A_{1},\dots,A_{r}$ of $A$ and compare for each $A_{i}$ the output of Mult$(X,Y)$ with Mult$(X,A_{i})$. If these sequences are equal we recursively go through the same procedure with $X$ and $A_{i}$. If they are not equal we append $A_{i}$ to a list of output components which we return in the end. This is potentially more optimal then minimizing after the stratification is finished because it prevents a build up of unneeded varieties.
@@ 4. Runtime Tests
Table 1. Run times of several examples, – denotes examples for which execution was stopped after 8 hours. These computations were run in version 1.24.5 of Macaulay2 on a workstation with an Intel Xeon W-3365 CPU and 1TB of RAM.

### v2

@@ 14 Return $W_{\bullet}$
The termination of the algorithm is clear. Note that the singular locus of any variety consists of the the union of the singular loci of its $\mathbb{Q}$-irreducible components together with the intersections of all its $\mathbb{Q}$-irreducible components. Thus lines 4 and 6 in Whitney guarantee the smoothness of all strata of the output of $\text{\bf Whitney}(X)$. Further, the fact that Whitney’s condition (B) is satisfied for all pair of the strata of the output of $\text{\bf Whitney}(X)$ follows immediately from Theorem˜3.6. ∎
Example 3.9.
We illustrate our algorithm Whitney by running it on the variety
$X:=\mathbf{V}(z(zx-y^{2}+tx^{3}))\subset\mathbb{C}^{4}.$
This is Example 4.1 in [14].
In the first iteration of the outermost for-loop starting on line 3 in $\text{\bf Whitney}(X)$ our algorithm decomposes $X$ into its irreducible components $Y_{1}:=\mathbf{V}(z)$ and $Y_{2}:=\mathbf{V}(zx-y^{2}+tx^{3})$. The irreducible intersection $Z:=\mathbf{V}(z,tx^{3}-y^{2})$ of $Y_{1}$ and $Y_{2}$ is added to $W_{2}$ in line 6 of Whitney. This intersection coincides with the singular locus of $X$, hence the next interesting step in our algorithm happens in line 11 when we call $\text{\bf WhitPoints}(Y_{1},Z)$. This call finds $h=x^{3}$ and hence the variety $Z\cap\mathbf{V}(x)=\mathbf{V}(x,y,z)$ is added to $W_{1}$ in line 13. Note that $\mathbf{V}(x,y,z)$ actually coincides with the singular locus of $Z$.
In the next iteration of the outermost for-loop in $\text{\bf Whitney}(X)$, we check Whitney’s condition (B) of the pairs consisting of $Z$ (which has dimension $2$) and any lower dimensional variety in our intermediate stratification $W_{\bullet}$, the only such lower dimensional variety is $\mathbf{V}(x,y,z)$. This turns out to be the last step where our stratification is updated, namely when we call $\text{\bf WhitPoints}(Z,\mathbf{V}(x,y,z))$ which returns $h=t$. Hence, the variety $\mathbf{V}(x,y,z,t)$ is added to $W_{\bullet}$. The output Whitney stratification finally consists of
$W_{3}=X=\mathbf{V}(z)\cup\mathbf{V}(zx-y^{2}+tx^{3})$
$W_{2}=Z=\mathbf{V}(z,tx^{3}-y^{2})$
$W_{1}=\mathbf{V}(x,y,z)$
$W_{0}=\mathbf{V}(x,y,z,t).$
@@
3.3. The Minimization Algorithm
Our first goal here is to show how to compute, for a given $X\subset\mathbb{C}^{n}$ and a random point $y$ in a subvariety $Y\subset X$, the sequence $m_{\bullet}(X,y)$ without being given $y$ explicitly. Next, we will show that the sequence $m_{\bullet}(X,\bullet)$ is constant on a Zariski-open subset of $Y$ if $Y$ is equidimensional. These two things combined allow us to compute, probabilistically, the “generic” multiplicities $m_{\bullet}(X,y)$, $y\in Y$. Combining this with Theorem˜2.15 we obtain a procedure to produce the unique minimal Whitney stratification of a complex algebraic variety given a Whitney stratification computed by Whitney.
For convenience we introduce the following notation: Given a sequence of polynomials $F:=(f_{1},\dots,f_{n})$ with $f_{i}\in\mathbb{C}[\mathbf{x}]$ we denote by $\phi_{F}$ the map
$\phi_{F}:\;$ $\mathbb{C}^{n}\rightarrow\mathbb{C}^{n}$
$p:=(p_{1},\dots,p_{n})\mapsto(f_{1}(p),\dots,f_{n}(p)).$
Recall that $\phi_{F}$ is said to be regular at a point $p\in\mathbb{C}^{n}$ if the Jacobian determinant of $F$ does not vanish at $p$.
Suppose for the moment that $Y$ is of codimension $c$ with $Y=\mathbf{V}(f_{1},\dots,f_{c})$ and that there are degree one polynomials $\ell_{c+1},\dots,\ell_{n}$ such that, letting $F:=(f_{1},\dots,f_{c},\ell_{c+1},\dots,\ell_{n})$, the map $\phi_{F}$ is regular at every point $y\in Y$.
[…]
@@ 3.3. The Minimization Algorithm
As $\phi_{F}$ was assumed to be regular at every $y\in Y$ it defines a local isomorphism near any $y\in Y$ by the inverse function theorem. This means that for any $y\in Y$, $\phi_{F}$ induces an isomorphism of completions
$\phi^{*}_{y,F}:\widehat{\mathscr{O}}_{\mathbb{C}^{n},\phi(y)}\rightarrow\widehat{\mathscr{O}}_{\mathbb{C}^{n},y}.$
[…]
We note that the proposition above is not surprising since it is known that polar multiplicities are analytic invariants of a germ of a reduced equidimensional complex analytic space $(Z,0)\subset(\mathbb{C}^{n},0)$ and are thus preserved by local isomorphism; see e.g. [4, Proposition 3.13] and the surrounding discussion and references.
Next suppose that $Y\subset X$ is equidimensional of codimension $c$ but no longer necessarily cut out by $f_{1},\dots,f_{c}$ as above. Then we can reduce to the case considered above as follows:
@@
Let $G$ be any generating set of the ideal $I(Y)$. For $c$ general linear combinations $f_{1},\dots,f_{c}$ of the elements in $G$ and general degree one polynomials $\ell_{1},\dots,\ell_{n-c}$ there is a Zariski-dense open subset $U$ of $Y$ such that the map $\phi_{F}$ is regular at every $y\in U$, where $F:=(f_{1},\dots,f_{c},\ell_{1},\dots,\ell_{n-c})$.
@@
For general $f_{1},\dots,f_{c}$ as in the statement of the proposition, $f_{1},\dots,f_{c}$ defines a maximal regular sequence in $I(Y)$. Hence there exists some polynomial $h$ such that $hI(Y)\subset\langle f_{1},\dots,f_{c}\rangle$. Hence the open subset $U_{1}:=Y-\mathbf{V}(h)$ is Zariski dense in $Y$ and, locally at every point $p\in U_{1}$, the ideal $I(Y)$ is generated by $f_{1},\dots,f_{c}$. To prove the proposition it therefore suffices to consider the case where $I(Y)=\langle f_{1},\dots,f_{c}\rangle$.
By applying Bertini’s theorem stated as in Theorem A.8.6 in [17] to every irreducible component of $Y$, we see that for for generic fixed degree one polynomials $\ell_{1},\dots,\ell_{n-c}$ there is a Zariski-open subset $U_{2}\subset\mathbb{C}^{n-c}$ such that for $\mathbf{a}:=(a_{1},\dots,a_{n-c})\in U_{2}$ the variety
$Z_{\mathbf{a}}:=\mathbf{V}(f_{1},\dots,f_{c},\ell_{1}-a_{1},\dots,\ell_{n-c}-a_{n-c})$
is zero-dimensional, does not meet the singular locus of $Y$ and is smooth. After potentially shrinking $U_{2}$, the variety $Z_{\mathbf{a}}$ intersects every irreducible component of $Y$ for every $\mathbf{a}\in U_{2}$. Let $F:=(f_{1},\dots,f_{c},\ell_{1},\dots,\ell_{n-c})$ and let $U$ be the preimage of $U_{2}$ under $\phi_{F}$ in $Y$. Note that $U$ is dense in $Y$. As $\langle f_{1},\dots,f_{c}\rangle$ is a radical ideal, the smoothness condition on $Z_{\mathbf{a}}$ means precisely that the Jacobian determinant of $F$ does not vanish at any point $p\in U$, proving the statement. ∎
@@
Let $m_{\bullet}(\tilde{X},0)$ be the multiplicity sequence constructed as at the beginning of this section, with $F$ chosen as in Proposition 3.12. Then $m_{\bullet}(\tilde{X},0)=m_{\bullet}(X,y)$ for every $y\in\phi^{-1}(0)$.
@@
Let $U\subset\mathbb{C}^{n}$ be as in Proposition 3.12. By the same proposition, with probability 11, the intersection $Y\cap\mathbf{V}(\ell_{c+1},\dots,\ell_{n})\cap U\subset\phi^{-1}(0)$ is not empty. For any choice of $y\in Y\cap\mathbf{V}(\ell_{c+1},\dots,\ell_{n})\cap U$ we then have $m_{\bullet}(\tilde{X},0)=m_{\bullet}(X,y)$ by Proposition 3.11. ∎
@@
5 Compute ideals defining $\phi(X)=\delta_{0}(\phi_{F}(X),0),\dots,\delta_{\dim(X)-1}(\phi_{F}(X),0)$
6 Using these ideals, Return $(m_{0}(\delta_{0}(\phi_{F}(X),0)),\dots,m_{0}(\delta_{\dim(X)-1}(\phi_{F}(X),0)))$
@@ 5 Compute ideals defining $\phi(X)=\delta_{0}(\phi_{F}(X),0),\dots,\delta_{\dim(X)-1}(\phi_{F}(X),0)$
In line 5, an ideal defining $\phi(X)$ is again computed using Gröbner basis computations with block orders, see e.g. Proposition 15.30 in [3]. Using the notation of Proposition 3.3, ideals defining the necessary local polar varieties can then be computed as ideals of the form
@@
In line 6, to compute the required multiplicities we use degree computations via Gröbner basis as described in [6] and implemented in the SegreClasses Macaulay2 package [5], see in particular [6, Theorem 5.3]. Alternatively one can use standard basis computations, see e.g. [16] for details.
Our next goal is the following: Suppose again that $Y\subset X$ is equidimensional. Then we want to show that the sequence $m_{\bullet}(X,\bullet)$ is constant on a Zariski-dense open subset of $Y_{\rm reg}$ and, since $Y_{\rm reg}$ is dense in $Y$, hence also on a Zariski-dense open subset of $Y$. This is concluded as follows from Theorem˜2.15:
@@
Given a variety $X$ and equidimensional $Y\subset X$, there exists a Zariski-dense open subset $U\subset Y_{\rm reg}$ such that $m_{\bullet}(X,\bullet)$ is constant on $U$.
@@
Suppose $\dim(Y)=e$. Using Proposition 3.12 and the same argument as in the proof of Proposition 3.11 we may suppose that $Y=\mathbf{V}(x_{e+1},\dots,x_{n})$ since we are computing local multiplicities and they are preserved by the map $\phi_{F}$ constructed in Proposition 3.11. Now, the set of points in $Y_{\rm reg}$ at which the pair $(X_{\rm reg},Y_{\rm reg})$ fails to satisfy Whitney’s condition (B) is contained in a proper Zariski-closed subset $W^{*}$ of $Y$, a suitable Zariski closed set $W^{*}$ is specified by Theorem 2.11 above, for example. Let $W$ be a Zariski closed set containing $Y_{\rm Sing}$ and $W^{*}$ and set U:=YWU:=Y-W. By Theorem˜2.15, every point $y\in U$ has a euclidean neighborhood on which the sequence $m_{\bullet}(X,\bullet)$ is equal to $m_{\bullet}(X,y)$. As $Y$ is just an affine subspace od $CC^{n}$, $U$ is connected, therefore $m_{\bullet}(X,\bullet)$ is constant on $U$, proving the theorem. ∎
@@ 7 Return $W_{\bullet}$
The termination is clear. Suppose that we have a triple of varieties $W\subset Z\subset Y$ with $W$ and $Z$ irreducible and that for generic $w\in W$ and $z\in Z$ we have $m_{\bullet}(Y,w)=m_{\bullet}(Y,z)$. Then there exists a Zariski-dense subset $U$ of $Z$ such that $U\cap W\neq\emptyset$ and such that $m_{\bullet}(Y,\bullet)$ is constant on $U$. Hence the points in $W$ at which the pair $(Y,Z)$ does not satisfy Whitney’s condition (B) have at least codimension one in $W$ by Theorem˜2.15. In the situation of WhitneyMinimize, $W$ can thus be removed without destroying the property of the output flag $W_{\bullet}$ being a Whitney stratification. This proves the correctness. ∎
@@
[…]
@@ 4. Runtime Tests
Note that in Table 1 we do not compare our implementation to the algorithms of [11] or of [13, 15] as all these algorithm are unable to successfully compute the Whitney stratification of the Whitney umbrella, the simplest non-trivial example, which is a surface in $\mathbb{C}^{3}$ defined by the binomial equation $x^{2}-y^{2}z=0$. In both cases we have let these algorithms run for greater than 24 hours on the Whitney umbrella on our test machine and neither finished. Additionally, in the case of the quantifier elimination based method of [13, 15], the Maple implementation of quantifier elimination eventually gives a stack limit exceeded error even if we allow it to make the stack arbitrarily large on our test machine (which has an Intel Xeon W-3365 CPU with 1032 GB of RAM). For comparison the Whitney umbrella example takes 0.1s to run with the algorithm of [7, 9] and 0.08s with the algorithm of Whitney algorithm of §3; both produce a minimal stratification.
@@
$X_{7}$, see (10) 15.4s No 20.2s
$X_{8}$, see (11) 1177.6s No 1934.6s
Table 1. Run times of several examples, – denotes examples for which execution was stopped after 24 hours. These computations were run in version 1.24.5 of Macaulay2 on a workstation with an Intel Xeon W-3365 CPU and 1TB of RAM.
@@
$\begin{split}X_{7}=\mathbf{V}(&x_{1}x_{2}^{3}-2x_{1}x_{2}^{2}x_{4}+x_{1}x_{2}x_{3}x_{5}+x_{1}x_{2}x_{4}^{2}+x_{1}x_{2}x_{5}-x_{1}x_{3}x_{4}x_{5}-x_{1}x_{4}x_{5}-x_{2}^{2}x_{3}x_{5}-\\ &x_{2}^{2}x_{4}-x_{2}^{2}x_{5}+x_{2}x_{3}x_{4}x_{5}-x_{2}x_{3}x_{5}+2x_{2}x_{4}^{2}+x_{2}x_{4}x_{5}-x_{2}x_{5}-x_{3}^{2}x_{5}^{2}+x_{3}x_{4}x_{5}-\\ &2x_{3}x_{5}^{2}-x_{4}^{3}+x_{4}x_{5}-x_{5}^{2})\end{split}$ (10)
$\begin{split}X_{8}=\mathbf{V}(&x_{1}^{2}x_{2}^{2}x_{3}-2x_{1}^{2}x_{2}x_{3}x_{4}-2x_{1}^{2}x_{2}x_{3}x_{5}+x_{1}^{2}x_{3}x_{4}^{2}+2x_{1}^{2}x_{3}x_{4}x_{5}+x_{1}^{2}x_{3}x_{5}^{2}-2x_{1}x_{2}^{3}x_{3}+\\ &4x_{1}x_{2}^{2}x_{3}x_{4}+4x_{1}x_{2}^{2}x_{3}x_{5}-x_{1}x_{2}^{2}x_{4}-2x_{1}x_{2}x_{3}x_{4}^{2}-6x_{1}x_{2}x_{3}x_{4}x_{5}-2x_{1}x_{2}x_{3}x_{5}^{2}+\\ &x_{1}x_{2}x_{4}^{2}+x_{1}x_{2}x_{4}x_{5}+2x_{1}x_{3}x_{4}^{2}x_{5}+2x_{1}x_{3}x_{4}x_{5}^{2}-x_{1}x_{4}^{2}x_{5}+x_{2}^{4}x_{3}-2x_{2}^{3}x_{3}x_{4}-\\ &2x_{2}^{3}x_{3}x_{5}+x_{2}^{3}x_{4}+x_{2}^{2}x_{3}x_{4}^{2}+4x_{2}^{2}x_{3}x_{4}x_{5}+x_{2}^{2}x_{3}x_{5}^{2}-2x_{2}^{2}x_{4}^{2}-x_{2}^{2}x_{4}x_{5}-2x_{2}x_{3}x_{4}^{2}x_{5}-\\ &2x_{2}x_{3}x_{4}x_{5}^{2}+x_{2}x_{4}^{3}+2x_{2}x_{4}^{2}x_{5}+x_{3}x_{4}^{2}x_{5}^{2}-x_{4}^{3}x_{5})\end{split}$ (11)